OpenStudy (anonymous):

A Ferris wheel with a diameter of 50 feet has 30 equally spaced passenger cars. What is the length of the arc between the centerline of two consecutive cars? Use the value π = 3.14, and round your answer to the nearest tenth of a foot.

OpenStudy (perl):

arclength = radius * theta (radians)

OpenStudy (perl):

so first find the angle between two consecutive cars , then plug into formula

OpenStudy (jdoe0001):

 \(\large \bf \textit{arc's length}=s=r\theta(radians)=\cfrac{r\theta(degrees) \pi}{180}\)